# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

### v1

@@ Abstract
Progress in enhancing large language model (LLM) planning and reasoning capabilities is significantly hampered by the bottleneck of scalable, reliable data generation and evaluation. To overcome this, I introduce NL2FLOW, a fully automated system for parametrically generating planning problems – expressed in natural language, a structured intermediate representation, and formal PDDL – and rigorously evaluating the quality of generated plans. I demonstrate NL2FLOW’s capabilities by generating a dataset of 2296 problems in the automated workflow generation domain and evaluating multiple open-sourced, instruct-tuned LLMs. My results reveal that the highest performing models achieved 86% success in generating valid plans and 69% in generating optimal plans, specifically for problems with feasible solutions. Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. Notably, I observed that the highest success rate for translating natural language into a JSON representation of a plan was lower than the highest rate of generating a valid plan directly. This suggests that unnecessarily decomposing the reasoning task – introducing intermediate translation steps – may actually degrade performance, implying a benefit to models capable of reasoning directly from natural language to action. As I scale LLM reasoning to increasingly complex problems, the bottlenecks and sources of error within these systems will inevitably shift. Therefore, a dynamic understanding of these limitations – and the tools to systematically reveal them – will be crucial for unlocking the full potential of LLMs as intelligent problem solvers.
@@ 1 Introduction
Figure 1: Steps to obtain a plan from a planning problem in natural language
Planning is crucial for modern conversational AI systems, enabling them to achieve user goals by leveraging available resources such as APIs, agents, tools, databases, and internet. While complex planning systems often rely on symbolic planners, building such systems is a significant undertaking. A key challenge lies in knowledge engineering – translating natural language problem descriptions into formal representations suitable for planning algorithms. This translation process is prone to error [1] and can become a bottleneck as system complexity increases. Furthermore, classical planning approaches, based on algorithms like those implemented in Planning Domain Definition Language (PDDL), can struggle with scalability.
@@
Classical AI planning, while offering guarantees like optimality, requires precise formalization of problems. Errors in translating natural language into PDDL can lead to planning failures. Large Language Models (LLMs) offer a promising alternative, demonstrating capabilities in language understanding, generation, and reasoning. However, LLM-based planning remains unreliable, and current approaches often focus on free-form plan generation or reinforcement learning within limited environments [2]. A significant advancement would be to leverage LLMs to reliably translate natural language planning problems into formal representations, or directly generate plans with high logical consistency (Figure 1). If an LLM can accurately translate or generate plans, the complexity of building a robust natural language planning system is drastically reduced.
However, progress in improving LLM planning and translation capabilities is hindered by a lack of scalable data generation and evaluation methods. To address this, I introduce NL2FLOW, a system designed to 1) generate planning problems expressed in natural language, an intermediate representation, and PDDL, and 2) evaluate the quality of generated plans. This study leverages NL2FLOW to investigate the inherent planning capabilities of open-source LLMs, establishing a baseline performance without task-specific optimization or architectural modifications. My primary goal is to identify domains where LLMs exhibit natural planning strengths and to develop a data generator capable of producing scalable, configurable planning datasets relevant to real-world applications. This will allow researchers to systematically probe the limitations of LLMs in complex planning scenarios, drawing on insights from automated planning to identify structural weaknesses in LLM-based approaches. Ultimately, this work aims to inform the development of improved architectures and training techniques for deep neural networks designed for planning tasks.
In this paper, I focus on applying LLMs to simple planning problems within the domain of workflow generation. Workflow creation – decomposing complex problems into structured, sequential tasks – is a common daily activity. I define “simple” problems as those solvable by symbolic planners within a short timeframe (e.g., a few seconds). I explore two primary approaches: 1) using LLMs to translate natural language problem descriptions into structured formats (e.g., JSON) suitable for symbolic planners, and 2) examining the direct generation of plans by LLMs from natural language descriptions.
This study is impactful in several key aspects:
  • Scalable Data Generation: NL2FLOW generates data randomly and parametrically, enabling scalable dataset creation. Scalability is primarily limited by the computational capacity of the symbolic planner used for data validation, ensuring the generated data is symbolically verifiable. The parametric nature of data generation also provides well-defined explanatory variables for downstream analysis. Critically, this approach requires no human or LLM intervention in data creation or inspection, making it ideal for training deep neural networks that require large datasets and programmatic reward computation.
  • Multi-Format Data: NL2FLOW data includes natural language descriptions, enabling the training and evaluation of LLMs for planning tasks, intermediate representations in python object, and PDDLs.
  • Symbolic Evaluation: The inclusion of PDDL allows for the symbolic evaluation of LLM-generated plans, providing mathematically guaranteed metrics like optimality. This enables scalable and reliable evaluation.
  • Problem Complexity Metric: I define problem complexity as the length of the optimal plan, simplifying downstream analysis and providing a clear measure of difficulty.
2 Related Work
Current evaluation of reasoning capacity often focuses on final answer accuracy for benchmarks like MATH500 [3], AIME24, GSM8K [4], and MATH. However, this metric overlooks the validity of the reasoning ’process’ itself; a correct answer can be achieved through flawed reasoning, effectively rewarding a successful, yet invalid, plan. Furthermore, relying solely on final answer correctness leaves the quality and structure of the reasoning process opaque. Similarly, BIRD [5], a natural language to SQL translation benchmark, primarily relies on final answer accuracy, despite the inherent planning aspects of SQL query construction alongside elements like intent detection. GSM-Symbolic [6] attempts to address this by generating mathematical problems from symbolic templates, allowing for manual verification of solutions. ReasonAgain [7] leverages GPT-4o to generate Python programs for GSM8K and MATH, validating reasoning by executing the generated code.
Beyond these benchmarks, several studies utilize “puzzle” datasets [8] - Tower of Hanoi, river crossing, blocks world, and checker jumping - to assess LLM reasoning. These puzzles offer the advantage of verifiable reasoning steps through simulation and can be scaled to challenge LLMs as they saturate. However, the applicability of these abstract puzzles to real-world scenarios remains limited.
A growing body of work focuses on evaluating and training Large Language Models (LLMs) for workflow automation and reasoning. Several datasets have been proposed to assess LLM planning capabilities, differing in their generation methods (static vs. dynamically generated) and evaluation approaches (manual vs. programmatic). This section details relevant work, highlighting these key distinctions.
2.1 Static Datasets & Primarily Manual/LLM-as-Judge Evaluation
FlowBench [9] presents a static dataset of 51 scenarios across 6 domains, focusing on multi-turn interactions. Evaluation relied heavily on human verification by comparing expected and observed traces, identifying a need for improved LLM planning. Similarly, HuggingGPT [10] utilizes a static, human-annotated dataset augmented with GPT-4 generated pseudo-labels, and employs LLM-as-a-judge for graph tasks. TaskLAMA [11] also provides a static dataset, created through human annotation, for structured complex task decomposition. ISG-BENCH [12] focuses on interleaved text and image generation, utilizing a static dataset and relying on MLLM-as-a-judge for scoring alignment, coherence, and helpfulness, grounded in human-annotated truth. WikiPlan and RecipePlan contribute static text-image pairs for task completion. These datasets, while valuable, often require significant human effort for creation and/or evaluation [13].
2.2 Static Datasets & Programmatic Evaluation
TaskBench [14] divides task automation into decomposition, tool selection, and parameter prediction, leveraging a static dataset of 17,331 samples generated with GPT-4. The authors employed programmatic metrics (F1, Normalized Edit Distance, Rouge) for evaluation, with some partial manual inspection. This approach offers scalability due to a large sample size but can be limited by the quality and coverage of the initial dataset. Open Grounded Planning [2] also aims to generate executable plans, providing LLMs with an extensive action library and leveraging existing datasets. Evaluation combines programmatic metrics (executability, pass rate) with LLM-as-a-judge for assessing completeness, feasibility, and relevance.
2.3 Algorithmically Generated Datasets & Programmatic Evaluation
My work builds upon these efforts by presenting a fully automated pipeline for both data generation and evaluation, addressing the limitations of static datasets and the need for scalable, programmatic (symbolic) assessment of LLM reasoning and planning capabilities.
@@ 3 Method
This section details the automated data generation and plan evaluation pipeline used in this study. I leveraged NL2FLOW (https://github.com/IBM/nl2flow) to generate a dataset of planning problems and evaluate plans produced by large language models.
@@ 3.1 Domain Description
This work explores automated planning within the domain of business process automation, specifically focusing on task flow generation. The system’s objective is to autonomously determine and execute a sequence of actions to achieve user-defined goals, such as submitting an expense report. I model each action as a function requiring a set of input parameters. Initially, the values of these parameters are often unknown, representing information the system needs to acquire. To address this, the system employs a dialogue-based approach. When a parameter value is missing, the system initiates a request for information via an ’ask’ action – a dialogue act designed to elicit the necessary value from the user. This process of identifying and requesting missing parameter values is known as slot-filling, a common technique in task-oriented dialogue systems. Successfully executing an action updates the system’s internal state, transitioning the corresponding parameter from an unknown to a known state. This allows the system to proceed with subsequent actions in the plan.
3.2 Planning Data Generation
Table 1: Planning data generator configuration
@@ 3.2 Planning Data Generation
Figure 2: Distribution of the number of steps in optimal plans across the dataset
I generated 2,296 planning data (intermediate representation + natural language descriptions + PDDL) using a parameterized approach (Table 1) to control the complexity and characteristics of the generated problems. Key parameters included the number of actions, action arity (the number of input and output parameters per action), the number of goal actions, the proportion of slot-fillable variables (variables requiring values to be determined during planning), and the proportion of coupled actions (actions where the output of one action serves as input to another). 30 variables were generated and designated as either slot-fillable or initially known. Actions were then created, establishing dependencies between them by connecting action outputs to subsequent action inputs. Variables not directly used within any action were assigned to the available data, whose values are already known. An action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the generated action set. The resulting intermediate representation (python object) was compiled into PDDL strings using NL2FLOW, facilitating symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication. The optimal plan length for the generated planning data ranged from 0 to 8 (Figure 2). A plan consisting of only one action was not feasible. Because the minimum arity of any action in this domain was 1, each action required at least one input variable. This input variable must be obtained either by invoking the ‘ask‘ action or another action, precluding a valid plan of length 1.
@@ 3.3 Prompt Generation
Listing 1: Verbose planning prompt
@@
Listing 2: Concise planning prompt
@@
Listing 3: Planning response example
@@
0. Execute action a__2. This will result in acquiring v__1. aa is required later by action A.
@@
2. Execute action a__3 with bb as input. Since executing a__3 was a goal of this plan, return the results of a__3(v__2) to the user.
@@
Listing 4: Translation prompt
@@
This study examined the capacity of large language models (LLMs) to translate natural language problem descriptions into JSON schemas and to generate executable plans to achieve specified goals. I adopted a few-shot learning approach, constructing prompt templates for both tasks and providing six example demonstrations. For the planning task, I investigated the impact of prompt length, specifically comparing ’verbose’ and ’desclarative’ prompt templates. The ’verbose’ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ’concise’ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt. For translation, a single prompt template (see Listing 4) was used, designed to elicit accurate and fluent translations. The complete prompt templates used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary materials to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area.
Each prompt template followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings 1 and 2). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing 3). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans. Example prompts are provided in the supplementary materials.

### v2

@@ Abstract
Effective agent performance relies on the ability to compose tools and agents into effective workflows. However, progress in Large Language Model (LLM) planning and reasoning is limited by the scarcity of scalable, reliable evaluation data. This study addresses this limitation by identifying a suitable workflow domain for LLM application. I introduce NL2Flow, a fully automated system for parametrically generating planning problems, which are expressed in natural language, a structured intermediate representation, and formal PDDL, and rigorously evaluating the quality of generated plans. NL2Flow generates a dataset of 2296 low-difficulty problems in automated workflow generation and evaluates multiple open-sourced, instruct-tuned LLMs without task-specific optimization or architectural modifications. Results reveal that the highest performing model achieved 86% success in generating valid plans and 69% in generating optimal plans, specifically for problems with feasible plans. Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. To investigate the potential of LLMs as natural language-to-JSON translators for workflow definition, and to facilitate integration with downstream symbolic computation tools and a symbolic planner, I evaluated the LLM’s translation performance on natural language workflow descriptions. I observed that translating natural language into a JSON representation of a workflow problem yielded a lower success rate than generating a plan directly, suggesting that unnecessary decomposition of the reasoning task may degrade performance and highlighting the benefit of models capable of reasoning directly from natural language to action. As LLM reasoning scales to increasingly complex problems, understanding the shifting bottlenecks and sources of error within these systems will be crucial.
@@ 1 Introduction
Automatic workflow generation is a critical capability for agentic systems. AI agents are now widely deployed to automate business processes across diverse industries, relying on tool use to access dynamic information, expand knowledge, and personalize interactions. In multi-agent systems, an orchestrating agent should compose agents to achieve goals. Effective agent performance hinges on the ability to efficiently ”select” and ”compose” tools or agents, essentially, to create effective workflows, in response to user requests. Traditional planning techniques, while conceptually sound, often fall short in addressing the scalability and adaptability requirements of these dynamic environments.
Figure 1: Steps to obtain a plan from a planning problem in natural language. Two approaches are shown: (1) Direct planning, where a Large Language Model (LLM) Planner directly generates a plan from the input. (2) Intermediate representation planning, where an LLM Translator first converts the natural language instruction into an intermediate representation, which is then used by symbolic planning algorithms to compute a plan.
Traditional AI planning systems rely on formal representations, such as those defined by Planning Domain Definition Language (PDDL). While these systems can offer guarantees like optimality, building them requires painstaking knowledge engineering to translate natural language goals into formal problem descriptions [1]. This translation process is a known bottleneck, and the complexity of maintaining and scaling these systems increases substantially with the number of tools and task complexities. For example, business automation systems can involve thousands of APIs. Consequently, symbolic planners struggle to find optimal workflows within reasonable timeframes, often requiring developers to manually reduce the scope of the problem by limiting the number of APIs considered.
@@
Large Language Models (LLMs) present a promising alternative, demonstrating impressive capabilities in natural language understanding, generation, and reasoning. However, leveraging LLMs for reliable planning remains challenging. Existing approaches often resort to free-form plan generation, which lacks logical consistency, or rely on reinforcement learning confined to limited environments [2]. A key advancement would be the ability to reliably translate natural language planning problems into formal representations or to directly generate logically sound and executable plans (Figure 1). Successful LLM-based planning would significantly reduce the engineering effort associated with building robust natural language planning systems.
Despite their potential, evaluating and improving LLM planning capabilities is hampered by the scarcity of scalable data generation and evaluation methodologies. This study addresses this need by introducing NL2Flow [3], a scalable data generation method for workflows. NL2Flow uses random and parametric approaches to create a dataset of natural language descriptions, Python objects, and PDDL representations, eliminating the need for human or LLM annotation and enabling the creation of large-scale datasets. I demonstrate that NL2Flow can generate workflows in a domain where open-sourced, instructed-tuned LLMs without task-specific optimization or architectural modifications exhibit promising planning performance for low-difficulty problems. Establishing a domain in which large language models (LLMs) exhibit promising reasoning and planning capabilities without task-specific optimization and modifications can enable researchers to investigate the origins of their domain-dependent reasoning/planning performance.
Importantly, the use of PDDL allows for symbolic evaluation of generated plans, providing mathematically grounded metrics, which include optimality, for reliable and scalable assessment. By systematically varying goals and agent/tool configurations, NL2Flow facilitates evaluation of LLMs both as planners and as translators from natural language to workflow representations, highlighting key bottlenecks in building robust, natural language-grounded AI planning systems (Figure 1). In this study, I focus on applying LLMs to low-difficulty workflow generation problems, which are readily solvable by symbolic planners, as a first step towards tackling more complex scenarios.
@@ 2 Method
This study used an automated planning integration system, NL2Flow [3], to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality.
@@ 2.1 Domain Description
Listing 1: NL2Flow domain example for booking a business trip [3]
[0] ask(conference name)
[1] ask(w3)
[2] list of papers = Author Workbench(w3, conference name)
[3] ask(username)
[4] invoice = Registration Bot(conference name, list of papers, username)
[5] map(invoice, conference registration)
[6] name, address, Employee ID, Passport = Travel Info Agent(w3)
[7] name, address, Employer Letter = Workday(Employee ID)
[8] visa = Visa Application(Passport, address, Employer Letter)
[9] ask(start date)
[10] ask(end date)
[11] map(home, address)
[12] map(BOS, destination)
[13] booking = Taxi(date, address, destination)
[14] map(LAX, address)
[15] map(JW Marriott Los Angeles LA 900 W Olympic Blvd, destination)
[16] booking = Taxi(date, address, destination)
[17] map(end date, date)
[18] map(JW Marriott Los Angeles LA 900 W Olympic Blvd, address)
[19] map(LAX, destination)
[20] booking = Taxi(date, address, destination)
[21] map(BOS, address)
[22] map(destination, home)
[23] booking = Taxi(date, address, destination)
[24] assert eval(is a business trip)
[25] flight_ticket, hotel_booking = Concur(start date, end date, home, destination)
[26] map(flight_ticket, ticket to conference)
[27] assert \$hotel_booking.price + \$flight_ticket.price < 1500
[28] approval = Trip Approval(ticket to conference, conference registration)
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows [4]. The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each custom action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a custom action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
2.2 Parametric Data Generation
Table 1: Workflow data generator configuration
@@ 2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of custom actions (representing available operations), custom action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. Custom actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. Custom actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A custom action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the custom action set. To simplify the problem and focus on core workflow generation challenges, variable mapping was excluded from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
2.2.1 Intermediate Representation
Listing 2: Intermediate representation from data generator. An example of the intermediate representation is provided in the supplementary material as ‘intermediate_representation.json‘.
available_agents”: [
agent_id”: a__2”,
actuator_signature”: {
in_sig_full”: [
name”: v__7”,
data_type”: null,
mappable”: false,
slot_fillable”: true
out_sig_full”: [
name”: v__9”,
data_type”: null,
mappable”: false,
slot_fillable”: true
goal_agent_ids”: [
a__1”,
a__0
mappings”: [],
available_data”: [
v__10”,
null
The parametric workflow problem generator constructed an intermediate representation (Listing 2), implemented as a Python object, to define the problem space. This representation comprises five key components:
  • $available\_agents$ represent the available custom actions
  • $actuator\_signature$ specifies the input and output parameters for each agent.
  • $mappings$ enable the transfer of values between variables.
  • $available\_data$ define the known values of variables at the workflow’s outset.
  • $goal\_agent\_ids$ are the ids of the desired actions to be achieved.
From this intermediate representation, I generated both natural language descriptions, using predefined templates, and a formal PDDL representation of the workflow problem.
@@ 2.3 Prompt Generation
Listing 3: Verbose planning prompt
@@
Listing 4: Concise planning prompt
@@
This study examined the capacity of LLMs to generate executable plans to achieve specified goals and to translate natural language problem descriptions into JSON schemas. I adopted a few-shot prompting approach, constructing prompt styles for both tasks and providing six example demonstrations. Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing 5). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans.
Listing 5: Planning response example
@@
0. Execute action a__2. This will result in acquiring v__1. v__1 is required later by action A.
@@
2. Execute action a__3 with v__2 as input. Since executing a__3 was a goal of this plan, return the results of a__3(v__2) to the user.
@@
For the planning task, I investigated the impact of prompt style, comparing ’verbose’ (Listing 3) and ’concise’ (Listing 4) prompt styles. The ’verbose’ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ’concise’ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings 6 and 7 show the description of custom action $a\_\_1$ in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing 6: Custom action in verbose prompt
Action a__1 has input parameters v__11, v__12, and v__13 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__11, v__12, and v__13 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Listing 7: Custom action in concise prompt
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__16, v__18
For translation, a single prompt style (Listing 8) was used, designed to elicit accurate and fluent translations. The complete prompt styles used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary material to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area. Each prompt style followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings 3 and 4). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Listing 8: Translation prompt
@@
This study evaluated nine LLMs across translation and planning tasks: CodeLlamma-34B-instruct, DeepSeek-coder-33B-instruct, DeepSeek-V3(685B, fp8), Granite-3.3-8B-instruct, Llama-3.1-8B-Instruct, Llama-3.1-405B-instruct-fp8, Llama-3.3-70B-instruct, Mixtral-8x22B-instruct-v0.1, and Qwen2.5-72B-instruct.